OpenStudy (anonymous):

What is the equation of the quadratic graph with a focus of (2, 0) and a directrix of y = -12?

hero (hero):

So you are given two points (2, 0) and (x, -12) Plug them into this formula: \((x - x_1)^2 + (y - y_1)^2 = (x - x_2)^2 + (y - y_2)^2\) Like so: \((x - 2)^2 + (y - 0)^2 = (x - x)^2 + (y - (-12))^2\) Then simplify

OpenStudy (anonymous):

oh! okay thank you so much!:)

hero (hero):

Let me know what you get after simplification.

OpenStudy (anonymous):

okay I never learned how to simplify something like this so I'm a bit stuck on the simplification

OpenStudy (anonymous):

I'm still trying though...

hero (hero):

Try your best...

hero (hero):

You have to expand the appropriate binomial squares first before you can simplify.

OpenStudy (anonymous):

Is that done by getting the square root of the equation? I have no clue what to do...

hero (hero):

Hints: \((x - 2)^2 = (x - 2)(x - 2)\) \(y - 0 = y\) \(x - x = 0\) \((y - (-12))^2 = (y + 12)^2 = (y + 12)(y + 12)\)

OpenStudy (anonymous):

Thing is I got y-0 and x-x 0 but I didn't think it'd work

OpenStudy (anonymous):

But I see how you did it.. makes sense

hero (hero):

Have you come up with a simplified form yet?

OpenStudy (anonymous):

I'm stuck on this now can I just get the square root of everything to eliminate the square so then I'll have x-2+y=y+12. So then I'll add 2 and subtract y to get x=14?

hero (hero):

No, you are a bit confused. You cannot take the square root of anything.

OpenStudy (anonymous):

Oh grr.. How do I do it then?

hero (hero):

You can take the square root of this: \((y - 2)^2 = (x + 2)^2\) But you cannot take the square root of this: \((x - 2)^2 + y^2 = (y +12)^2\)

hero (hero):

You must expand the binomial squares first. I gave you a hint on how to do that.

OpenStudy (anonymous):

Alright I'll try now

OpenStudy (anonymous):

In this equation though (x−2)^2 +y^2 =(y+12)^2 Can't you just have x=14 and y=0 though.

OpenStudy (anonymous):

It left me wondering if that could be it

hero (hero):

Do you know how to multiply (x - 2)(x - 2) ?

OpenStudy (anonymous):

Distributing..?

hero (hero):

And how would you do that?